The base of a triangle is 5 cm greater than the height .the area is 18cm^2 find height and length of base?
Answer by scott8148(6628)

let x=height, so x+5=base __ area=(1/2)base*height __ 18=(1/2)(x+5)(x)

multiplying by 2 __ 36=x^2+5x

subtracting 36 __ 0=x^2+5x-36

factoring __ 0=(x+9)(x-4)